# DySLIM: Dynamics Stable Learning by Invariant Measure for Chaotic Systems

### v1

@@ Abstract
Learning dynamics from dissipative chaotic systems is notoriously difficult due to their inherent instability, as formalized by their positive Lyapunov exponents, which exponentially amplify errors in the learned dynamics. However, many of these systems exhibit ergodicity and an attractor: a compact and highly complex manifold, to which trajectories converge in finite-time, that supports an invariant measure, i.e., a probability distribution that is invariant under the action of the dynamics, which dictates the long-term statistical behavior of the system. In this work, we leverage this structure to propose a new framework that targets learning the invariant measure as well as the dynamics, in contrast with typical methods that only target the misfit between trajectories, which often leads to divergence as the trajectories’ length increases. We use our framework to propose a tractable and sample efficient objective that can be used with any existing learning objectives. Our Dynamics Stable Learning by Invariant Measures (DySLIM) objective enables model training that achieves better point-wise tracking and long-term statistical accuracy relative to other learning objectives. By targeting the distribution with a scalable regularization term, we hope that this approach can be extended to more complex systems exhibiting slowly-variant distributions, such as weather and climate models.
Yair Schiff
Department of Computer Science
Cornell Tech
New York, NY, USA
yairschiff@cs.cornell.edu
&Zhong Yi Wan
Google Research
Moutain View, CA
Santa Narimana, Levand
stariate@ee.mount-sheikh.edu
@@ 1 Introduction
Building data-driven surrogate models to emulate the dynamics of complex time-dependent systems is one of the cornerstone tasks in scientific machine learning (Farmer & Sidorowich, 1987), with applications ranging from fluid dynamics (Sanchez-Gonzalez et al., 2020), weather forecasting (Pathak et al., 2022; Bi et al., 2023; Lam et al., 2022), climate modeling (Kochkov et al., 2023), molecular dynamics (Jia et al., 2020; Merchant et al., 2023), quantum chemistry (Chen et al., 2020; Zepeda-Núñez et al., 2021), and plasma physics (Anirudh et al., 2022; Mathews et al., 2021).
@@
Historically, various methods based on PCA (Pearson, 1901) and Koopman theory (Koopman, 1931) have been proposed to learn such emulators by leveraging large datasets to build a surrogate model during a, typically expensive, offline phase (Alexander & Giannakis, 2020; Schmid, 2010, 2022; Kaiser et al., 2021). The learned emulator provides fast and inexpensive inference, which is then used to accelerate downstream tasks such as design, control, optimization, data assimilation, and uncertainty quantification. Alas, many of these techniques are inherently linear, which renders them inadequate for problems with highly non-linear dynamics.
@@
Indeed, many of the underlying physical processes driving target applications are described by non-linear and chaotic systems, which are characterized by strong instabilities, particularly with respect to initial conditions: trajectories with close initial conditions diverge quickly due to the positive Lyapunov exponents (Medio & Lines, 2001; Strogatz, 2018). Fortunately, many of these system are dissipative, which implies the existence of a compact set, often called an attractor, towards which all bounded sets of initial conditions converge in finite time (Temam, 2012). In addition, many of these systems also empirically exhibit ergodicity, whose main consequence translates to the existence of an attractor-induced invariant measure, a measure that is invariant to the dynamics of the system, which captures the long-term behavior of the system (Stuart & Humphries, 1998).
Refer to caption (b) Refer to caption (c)
@@
Figure 1: Improved stability with regularized objectives in the KS and Lorenz 63 systems. (a) Sample ground truth and predicted trajectory across models trained on the KS system using Curriculum training (Curr) and the Pushforward trick (Pfwd) with/without regularization. The base versions showcase the blow-up issue (Curr) and wrong long-time dynamics (Pfwd). (b) Sinkhorn Divergence (SD)↓ between trajectories at various rollout times of the Lorenz $63$ system. Each point represents a random training seed, with the solid line indicating median values. (c) Cosine similarity↑ over time for the Lorenz $63$ system. Each line corresponds to one of five random training seeds with bolded lines indicating median values.
@@
Recent advances in machine learning (ML), have driven the development of several techniques for learning data-driven surrogates for non-linear dynamics (Rajendra & Brahmajirao, 2020; Roy & Rana, 2021; Brunton & Kutz, 2022; Ghadami & Epureanu, 2022; Nghiem et al., 2023). In the context of data-driven learning, autoregressive models (the focus of this paper) are a prevalent approach due to their ability to infer trajectories of arbitrary length. These autoregressive models predict a system’s state at time $t+\Delta t$ based on its state at time $t$. Iterative application (unrolling) allows for trajectory forecasting far beyond training time horizons.
@@
However, learning chaotic dynamics using autoregressive models in a stable manner has proven elusive. Due to memory and computational constraints, traditional ML-based approaches focus on learning short-term dynamics by minimizing the mean square error (MSE) between reference trajectories and those generated by unrolling a learned model; commonly using recurrent neural networks (Vlachas et al., 2018; Fan et al., 2020) or learning a projection from a stochastic trajectory using reservoir computing (Pathak et al., 2017; Bollt, 2021; Hara & Kokubu, 2022). Unfortunately, these models usually overfit to the short-term dynamics to the detriment of accurately predicting long-term behavior (Bonev et al., 2023). This manifests as trajectory blow-up, the values of the state variables diverging to infinity, or inaccurate long-term statistics during long-time inference as depicted in Figure 1 (a). Recent works have focused on minimizing the misfit between increasingly longer trajectories (Keisler, 2022). Although these methods have shown to attenuate the instability, the underlying difficulty remains: due to chaotic divergence, the losses become increasingly uninformative, which causes their gradients [^1] Which are computed by backpropagation through the unrolled steps and are prone to exacerbate instabilities in the system. This is related to the well-known issue of exploding/vanishing gradients (Pascanu et al., 2013). to diverge, as shown by Mikhaeil et al. (2022).
@@
A prime example of this challenge is weather and climate. While state-of-the-art ML models can learn short-term weather patterns (Lam et al., 2022), learning long-term climate behavior remains a very challenging open problem (Kochkov et al., 2023; Watt-Meyer et al., 2023). Thus the question we seek to answer becomes: How do we incorporate knowledge of a system’s long-term behavior into the learning stage, so that the model remains point-wise accurate for short-term predictions and statistically accurate for long-term ones.
@@
This work addresses this challenge by leveraging the presence of attractors and their invariant measures. We propose a framework that directly targets the learning of long-term statistics through sampling from the invariant measure induced by the ML model, enforced by a measure-matching regularization loss.
Contributions
@@
The contributions of this paper are three-fold: first, we propose a probabilistic and scalable framework for learning chaotic dynamics using data-driven, ML-based methods. Our framework introduces a system-agnostic measure-matching regularization term into the loss that induces stable and accurate trajectories satisfying the long-term statistics while enhancing short-term predictive power. Second, we use our framework to propose a tractable and sample-efficient objective that we dub Dynamics Stable Learning by Invariant Measures (DySLIM) that can be used in conjunction with any existing dynamical system learning objective. Third, we demonstrate that DySLIM is capable of tackling larger and more complex systems than competing probabilistic methods, up to a state-dimension of 4,096 with complex 2D dynamics. Namely, we show competitive results in three increasingly complex and higher dimensional problems: the Lorenz $63$ system (Lorenz, 1963; Tucker, 1999), a prototypical chaotic systems with the well known “butterfly” attractor, the Kuramoto-Sivashinsky (KS) equation, a 1D chaotic PDE, and the Kolmogorov-Flow (Obukhov, 1983), a 2D chaotic system that is routinely used as a benchmark (Kochkov et al., 2021) for turbulent fluid dynamics. For the largest systems, we show that DySLIM performance remains stable even for large batch sizes and learning rates, regimes in which the performance of other methods deteriorates rapidly. These capabilities are potentially useful for accelerating the training stage by leveraging data parallelism and large learning rates.
@@ 2 Background
where ${\mathcal{S}}_{\Delta t}$ is the map obtained by integrating Equation 1 in time by a period $\Delta t$. In what follows, for brevity, we drop the subscript $\Delta t.$ Then we can use the operator ${\mathcal{S}}$ to unroll, or advance in time, the solution of the dynamical system,
@@
Chaotic Systems
A chaotic system can be loosely defined as one whose trajectories are highly sensitive to initial conditions. Let $({\mathcal{U}},\mathrm{d})$ be a Euclidean metric space. We say that the system given by ${\mathcal{S}}$ is chaotic if there exists $\varepsilon>0$ such that for all ${\bm{u}}_{0}\in{\mathcal{U}}$ and $\delta>0,$ there exists ${\bm{v}}_{0}\in B_{\delta}({\bm{u}}_{0})$ and $k\in{\mathbb{N}},$ such that:
@@
Invariant Measures and Attractors
We assume that the state space is measurable $({\mathcal{U}},{\mathcal{A}}),$ where ${\mathcal{A}}$ is the Borel $\sigma$-algebra on $({\mathcal{U}},\mathrm{d}),$ and we have a probability measure $\mu:{\mathcal{A}}\rightarrow[0,1].$ If the discrete time dynamical system map ${\mathcal{S}}$ is measurable, then it also defines a probability distribution ${\mathcal{S}}_{\#}\mu:{\mathcal{A}}\rightarrow[0,1]$ which is called the pushforward of $\mu$ by ${\mathcal{S}}$, with ${\mathcal{S}}_{\#}\mu(A)=\mu({\mathcal{S}}^{-1}(A)),$ for all $A\in{\mathcal{A}}.$ We say ${\mathcal{S}}$ preserves a measure $\mu$, also denoted as $\mu$ is invariant [^2] We note that invariant measures may not be unique. For example, transformations with high degree of symmetries, such as a rigid-body transformation (e.g., translation and rotation), can have an infinite number of invariant measures. under/to ${\mathcal{S}},$ if:
@@ 3 Learning Dynamical Systems
Our goal is to find a parametric model ${\mathcal{S}}_{\theta}$ that governs our system in a manner consistent with the true dynamics defined by ${\mathcal{S}}$. To do so, we leverage previously collected data, which consists of $n$ trajectories: ${\mathcal{D}}=\{({\bm{u}}^{(i)}_{j})_{j=0}^{\ell^{(i)}}\}_{i=1}^{n},$ where $\ell^{(i)}$ is the length of the $i$-th trajectory, whose initial conditions are sampled from invariant measure supported on the attractor, i.e., $\{{\bm{u}}^{(i)}_{0}\}_{i=1}^{n}\stackrel{{\scriptstyle iid}}{{\sim}}\mu_{0}=\mu^{*}$. Letting $\mu_{j}$ be the distribution over states ${\bm{u}}_{j}$, i.e., states after $j$ time-steps, for ergodic dissipative systems, we have that $\mu_{j}:={\mathcal{S}}^{j}_{\sharp}\mu_{0}={\mathcal{S}}^{j}_{\sharp}\mu^{*}=\mu^{*}$.
@@
At inference, the learned models generate trajectories by autoregressively unrolling predictions, as in Equation 3: starting from a given ${\bm{u}}_{0},$ we generate $\tilde{{\bm{u}}}_{k}={\mathcal{S}}_{\theta}^{k}({\bm{u}}_{0}).$ As we unroll for large $k$ the learning dynamics can become unstable, by either diverging or converging to a different attractor.
@@
To attenuate this issue, two popular objectives have been introduced recently, which have been used to train state-of-the-art models (Brandstetter et al., 2022; Kochkov et al., 2023; Lam et al., 2022). Specifically, we examine a generalization of ${\mathcal{L}}^{\text{1-step}}$, the $\ell$-step objective:
@@
where $\omega(k)$ is a discount factor used to stabilize the training [^3] Since matching further rolled-out steps increases in difficulty with $k,$ especially for chaotic systems, we consider a monotonically decreasing discount factor of the form $\omega(k)=r^{k-1},$ $10<r<$ 1 inspired by Kochkov et al. (2023).. Training paradigms where $\ell$ starts at one and is gradually increased are known as curriculum training (Curr; Keisler (2022)), and we denote them as ${\mathcal{L}}^{\text{Curr}}$.
@@
${\mathcal{L}}^{\text{Pfwd},\ell}(\theta)=\mathbb{E}_{j}\mathbb{E}_{{\bm{u}}_{j}\sim\mu_{j}}[\omega(\ell)||{\mathcal{S}}_{\theta}(\mathtt{sg}({\mathcal{S}}^{\ell-1}_{\theta}({\bm{u}}_{j})))-{\mathcal{S}}^{\ell}({\bm{u}}_{j})||^{2}].$ (7)
This objective can either be used to replace or in addition to those defined Equations 5 and 6.
Refer to caption (a) Refer to caption (b)
Figure 2: Regularized objectives outperform baselines for the KS system. (a) Cosine similarity over time. Each line corresponds to the mean over trajectories of each of five random training seeds, with bold lines indicating median values. (b) Sinkhorn Divergence (SD) between trajectories at various rollout times. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Sources of Instability
We recast the instability of learned dynamical models as short-term overfitting and long-term distribution shift: parameters $\theta$ that minimize ${\mathcal{L}}^{\text{1-step}}$ on training data often overfit to this data and lead to ${\mathcal{S}}_{\theta\#}\mu_{j}\neq\mu_{j+1}.$ When deployed, the learned dynamical model will accumulate errors along a predicted trajectory as the distribution of predicted states veers further away from that of the actual system. Recent techniques (including the Curr and Pfwd training) attempt to mitigate this issue by encouraging the model to recover from deviations caused by pushing forward by ${\mathcal{S}}_{\theta}$. However, these objectives are far from perfect. For example, Figure 1 (a) depicts issues for the chaotic KS equation. The model trained with Curr training fails to generalized beyond the first $\ell$ steps in which was trained: the trajectory quickly enters an unstable attractor, from which it blows up. Similarly, the model trained using the Pfwd training, is able to learn the short-term dynamics, however, as time increases the trajectories enter a different attractor, one in which the dynamics are biased towards the right. In both cases, by introducing our proposed regularization, we are able to correct the long-term behavior.

### v2

@@ Abstract
Learning dynamics from dissipative chaotic systems is notoriously difficult due to their inherent instability, as formalized by their positive Lyapunov exponents, which exponentially amplify errors in the learned dynamics. However, many of these systems exhibit ergodicity and an attractor: a compact and highly complex manifold, to which trajectories converge in finite-time, that supports an invariant measure, i.e., a probability distribution that is invariant under the action of the dynamics, which dictates the long-term statistical behavior of the system. In this work, we leverage this structure to propose a new framework that targets learning the invariant measure as well as the dynamics, in contrast with typical methods that only target the misfit between trajectories, which often leads to divergence as the trajectories’ length increases. We use our framework to propose a tractable and sample efficient objective that can be used with any existing learning objectives. Our Dynamics Stable Learning by Invariant Measure (DySLIM) objective enables model training that achieves better point-wise tracking and long-term statistical accuracy relative to other learning objectives. By targeting the distribution with a scalable regularization term, we hope that this approach can be extended to more complex systems exhibiting slowly-variant distributions, such as weather and climate models. Code to reproduce our experiments is available here.
Machine Learning, ICML
@@ 1 Introduction
Building data-driven surrogate models to emulate the dynamics of complex time-dependent systems is a cornerstone task in scientific machine learning (Farmer & Sidorowich, 1987), with applications ranging from fluid dynamics (Sanchez-Gonzalez et al., 2020), weather forecasting (Lam et al., 2022; Pathak et al., 2022; Bi et al., 2023), climate modeling (Kochkov et al., 2023), molecular dynamics (Jia et al., 2020; Merchant et al., 2023), quantum chemistry (Chen et al., 2020; Zepeda-Núñez et al., 2021), and plasma physics (Mathews et al., 2021; Anirudh et al., 2022).
@@
Historically, various methods based on PCA (Pearson, 1901) and Koopman theory (Koopman, 1931) have been proposed to learn emulators by leveraging large datasets to build a surrogate model during a, typically expensive, offline phase (Schmid, 2010; Alexander & Giannakis, 2020; Kaiser et al., 2021; Schmid, 2022). The learned emulator provides fast and inexpensive inference, which is then used to accelerate downstream tasks such as design, control, optimization, data assimilation, and uncertainty quantification. Alas, many of these techniques are inherently linear, which renders them inadequate for problems with highly non-linear dynamics.
@@
Indeed, many of the underlying physical processes driving target applications are described by non-linear and chaotic systems, which are characterized by strong instabilities, particularly with respect to initial conditions: trajectories with close initial conditions diverge quickly due to the positive Lyapunov exponents (Medio & Lines, 2001; Strogatz, 2018). Fortunately, many of these systems are dissipative, which implies the existence of a compact set, often called an attractor, towards which all bounded sets of initial conditions converge in finite-time (Temam, 2012). In addition, many of these systems also empirically exhibit ergodicity, whose main consequence translates to the existence of an attractor-induced invariant measure, a measure that is unchanged by the dynamics of the system, which captures the long-term behavior of the system (Stuart & Humphries, 1998).
@@
Figure 1: Improved stability with regularized DySLIM objectives in the Kuramoto–Sivashinsky (KS) and Lorenz 63 systems. (a) Sample ground truth and predicted trajectory across models trained on the KS system using Curriculum training (Curr) and the Pushforward trick (Pfwd) with/without regularization. The base versions showcase the blow-up issue (Curr) and wrong long-time dynamics (Pfwd). (b) Sinkhorn Divergence (SD; $\downarrow$) between trajectories at various rollout times of the Lorenz 63 system. Each point represents a random training seed, with the solid line indicating median values. (c) Cosine similarity ($\uparrow$) over time for the Lorenz 63 system. Each line corresponds to one of five random training seeds with bolded lines indicating median values.
@@
Recent advances in machine learning (ML) have driven the development of several techniques for learning data-driven surrogates for non-linear dynamics (Rajendra & Brahmajirao, 2020; Roy & Rana, 2021; Brunton & Kutz, 2022; Ghadami & Epureanu, 2022; Nghiem et al., 2023). In the context of data-driven learning, autoregressive models (the focus of this paper) are a prevalent approach due to their ability to infer trajectories of arbitrary length. These autoregressive models predict a system’s state at time $t+\Delta t$ based on its state at time $t$. Iterative application (unrolling) allows for trajectory forecasting far beyond training time horizons.
@@
However, learning chaotic dynamics using autoregressive models in a stable manner has proven elusive. Due to memory and computational constraints, traditional ML-based approaches focus on learning short-term dynamics by minimizing the mean square error (MSE) between reference trajectories and those generated by unrolling a learned model; commonly using recurrent neural networks (Vlachas et al., 2018; Fan et al., 2020) or learning a projection from a stochastic trajectory using reservoir computing (Pathak et al., 2017; Bollt, 2021; Hara & Kokubu, 2022). Unfortunately, these models usually overfit to the short-term dynamics to the detriment of accurately predicting long-term behavior (Bonev et al., 2023). This manifests as trajectory blow-up, the values of the state variables diverging to infinity, or inaccurate long-term statistics during inference with large time-horizon, as depicted in Figure 1 (a). Recent works have focused on minimizing the misfit between increasingly longer trajectories (Keisler, 2022). Although these methods have been shown to attenuate the instability, the underlying difficulty remains: due to chaotic divergence, the losses become increasingly uninformative, which causes their gradients [^1] Gradients are computed by backpropagation through the unrolled steps and are prone to exacerbate instabilities in the system. This is related to the well-known issue of exploding/vanishing gradients (Pascanu et al., 2013). to diverge, as shown by Mikhaeil et al. (2022).
@@
A prime example of this challenge is weather and climate. While state-of-the-art ML models can learn short-term weather patterns (Lam et al., 2022), learning long-term climate behavior remains a very challenging open problem (Kochkov et al., 2023; Watt-Meyer et al., 2023). Thus the question we seek to answer becomes: How do we incorporate knowledge of a system’s long-term behavior into the learning stage, so that models remain point-wise accurate for short-term predictions and statistically accurate for long-term ones.
@@
This work addresses this challenge by leveraging the presence of attractors and their invariant measures. We propose a framework that directly targets the learning of long-term statistics by a measure-matching regularization loss.
@@
Contributions   The contributions of this paper are three-fold: first, we propose a probabilistic and scalable framework for learning chaotic dynamics using data-driven, ML-based methods. Our framework introduces a system-agnostic [^2] We assume no knowledge of the systems, such as the expression of the equations driving the dynamics. measure-matching regularization term into the loss that induces stable and accurate trajectories satisfying the long-term statistics while enhancing short-term predictive power. Second, we use our framework to propose a tractable, sample and computationally-efficient [^3] Our regularization loss incurs an extra cost depending only quadratically on the batch size. objective that we dub Dynamics Stable Learning by Invariant Measure (DySLIM) that can be used in conjunction with any existing dynamical system learning objective. Third, we demonstrate that DySLIM is capable of tackling larger and more complex systems than competing probabilistic methods, up to a state-dimension of 4,096 with complex 2D dynamics. Namely, we show competitive results in three increasingly complex and higher dimensional problems: the Lorenz 63 system (Lorenz, 1963; Tucker, 1999), a prototypical chaotic systems with the well known “butterfly” attractor, the Kuramoto-Sivashinsky (KS) equation, a 1D chaotic PDE, and the Kolmogorov-Flow (Obukhov, 1983), a 2D chaotic system that is routinely used as a benchmark for turbulent fluid dynamics (Kochkov et al., 2021). For the largest systems, we show that DySLIM performance remains stable even for large batch sizes and learning rates, regimes in which the performance of other methods deteriorates rapidly. These capabilities are potentially useful for accelerating the training stage by leveraging data parallelism and large learning rates.
@@ 2 Background
where ${\mathcal{S}}_{\Delta t}$ is the map obtained by integrating Equation 1 in time by a period $\Delta t$. In what follows, for brevity, we drop the subscript $\Delta t.$ We can use the operator ${\mathcal{S}}$ to unroll, or advance in time, the solution of the dynamical system,
@@
Chaotic Systems   A chaotic system can be loosely defined as one whose trajectories are highly sensitive to initial conditions. Let $({\mathcal{U}},\mathrm{d})$ be an Euclidean metric space. We say that the system given by ${\mathcal{S}}$ is chaotic if there exists $\varepsilon>0$ such that for all ${\bm{u}}_{0}\in{\mathcal{U}}$ and $\delta>0,$ there exists ${\bm{v}}_{0}\in B_{\delta}({\bm{u}}_{0})$ and $k\in{\mathbb{N}},$ such that:
@@
Invariant Measures and Attractors   We assume that the state space is measurable $({\mathcal{U}},{\mathcal{A}}),$ where ${\mathcal{A}}$ is the Borel $\sigma$-algebra on $({\mathcal{U}},\mathrm{d}),$ and we have a probability measure $\mu:{\mathcal{A}}\rightarrow[0,1].$ If the discrete-time dynamical system map ${\mathcal{S}}$ is measurable, then it also defines a probability distribution ${\mathcal{S}}_{\#}\mu:{\mathcal{A}}\rightarrow[0,1]$ which is called the pushforward of $\mu$ by ${\mathcal{S}}$, with ${\mathcal{S}}_{\#}\mu(A)=\mu({\mathcal{S}}^{-1}(A)),$ for all $A\in{\mathcal{A}}.$ We say ${\mathcal{S}}$ preserves a measure $\mu$, also denoted as $\mu$ is invariant [^4] We note that invariant measures may not be unique. For example, transformations with high degree of symmetries, such as a rigid-body transformation (e.g., translation and rotation), can have an infinite number of invariant measures. under/to ${\mathcal{S}},$ if:
@@ 3 Learning Dynamical Systems
Our goal is to find a Markovian parametric model ${\mathcal{S}}_{\theta}$ that governs our system in a manner consistent with the true dynamics defined by ${\mathcal{S}}$. To do so, we leverage previously collected data, which consists of $n$ trajectories: ${\mathcal{D}}=\{({\bm{u}}^{(i)}_{j})_{j=0}^{\ell^{(i)}}\}_{i=1}^{n},$ where $\ell^{(i)}$ is the length of the $i$-th trajectory, whose initial conditions are sampled from an invariant measure supported on the attractor, i.e., $\{{\bm{u}}^{(i)}_{0}\}_{i=1}^{n}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mu_{0}=\mu^{*}$. Letting $\mu_{j}$ be the distribution over states ${\bm{u}}_{j}$, i.e., states after $j$ time-steps, for ergodic dissipative systems, we have that $\mu_{j}:={\mathcal{S}}^{j}_{\sharp}\mu_{0}={\mathcal{S}}^{j}_{\sharp}\mu^{*}=\mu^{*}$.
@@
At inference, learned models generate trajectories by autoregressively unrolling predictions, as in Equation 3: starting from a given ${\bm{u}}_{0},$ we generate $\tilde{{\bm{u}}}_{k}={\mathcal{S}}_{\theta}^{k}({\bm{u}}_{0}).$ As we unroll for large $k$, the learning dynamics can become unstable, by either diverging or converging to a different attractor.
@@
To attenuate this issue, two popular objectives have been introduced recently, which have been used to train state-of-the-art models (Brandstetter et al., 2022; Lam et al., 2022; Kochkov et al., 2023). Specifically, we examine a generalization of ${\mathcal{L}}^{\text{1-step}}$, the $\ell$-step objective:
@@
where $\omega(k)$ is a discount factor used to stabilize training [^5] Since matching further rolled-out steps increases in difficulty with $k,$ especially for chaotic systems, we consider a monotonically decreasing discount factor of the form $\omega(k)=r^{k-1},$ $10<r<$ 1, inspired by Kochkov et al. (2023).. Training paradigms where $\ell$ starts at one and is gradually increased are known as curriculum training (Curr; Krishnapriyan et al. (2021); Keisler (2022)), and we denote them as ${\mathcal{L}}^{\text{Curr}}$.
@@
${\mathcal{L}}^{\text{Pfwd},\ell}(\theta)=$ (7)
$\mathbb{E}_{j}\mathbb{E}_{{\bm{u}}_{j}\sim\mu_{j}}[\omega(\ell)||{\mathcal{S}}_{\theta}(\mathtt{sg}({\mathcal{S}}^{\ell-1}_{\theta}({\bm{u}}_{j})))-{\mathcal{S}}^{\ell}({\bm{u}}_{j})||^{2}].$
This objective can either be used to replace or in addition to those defined in Equations 5 and 6.
Figure 2: Regularized DySLIM objectives outperform baselines for the KS system. (a) Cosine similarity ($\uparrow$) over time. Each line corresponds to the mean over trajectories of each of five random training seeds, with bold lines indicating median values. (b) Sinkhorn Divergence (SD; $\downarrow$) between trajectories at various rollout times. Each point represents a random training seed that remains stable, with the solid line indicating median values.
Sources of Instability   We recast the instability of learned dynamical models as short-term overfitting and long-term distribution shift: parameters $\theta$ that minimize ${\mathcal{L}}^{\text{1-step}}$ on training data often overfit to this data and lead to ${\mathcal{S}}_{\theta\#}\mu_{j}\neq\mu_{j+1}.$ When deployed, the learned dynamical model will accumulate errors along a predicted trajectory as the distribution of predicted states veers further away from that of the actual system. Recent techniques (including the Curr and Pfwd training) attempt to mitigate this issue by encouraging the model to recover from deviations caused by pushing forward by ${\mathcal{S}}_{\theta}$. However, these objectives are still prone to instabilities. For example, Figure 1 (a) depicts issues for the chaotic KS equation. The model trained with Curr training fails to generalize beyond the first $\ell$ steps on which it was trained: the trajectory quickly enters an unstable attractor, from which it blows up. Similarly, the model trained using the Pfwd training is able to learn the short-term dynamics, however, as time increases, the trajectories enter a different attractor, one in which the dynamics are biased towards the right. In both cases, by introducing our proposed regularization, we are able to correct the long-term behavior.